OpenStudy (anonymous):

solve the equation 2 over 3 (x − 4) = 2x for x. 2 −2 −8 −4

OpenStudy (blurbendy):

\[\frac{ 2 }{ 3(x-4) } = 2x?\]

OpenStudy (anonymous):

multiply both sides by \(\frac{3}{2}\) first

OpenStudy (anonymous):

\[\frac{2}{3}(x-4)=2x\] \[x-4=\frac{2}{2}\times 2x=32\] solve \[x-4=3x\] in two steps

OpenStudy (anonymous):

typo, i meant \[x-4=\frac{3}{2}\times 2x=32\]

OpenStudy (anonymous):

ugh im so confused

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

the question is \[\frac{2}{3}(x-4)=2x\] right?

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

ok so to get rid of that \(\frac{2}{3}\) out front of the parentheses, multiply both sides by the reciprocal, \(\frac{3}{2}\)

OpenStudy (anonymous):

that just leaves \(x-4\) on the left hand side of the equal sign the right hand side will be \(\frac{3}{2}\times 2x=3x\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait did you divide? I suck at this haha

OpenStudy (anonymous):

i multiplied \(\frac{3}{2}\times 2x=\frac{3}{\cancel{2}}\cancel{2}x=3x\)

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

in other words, three halves of two is three

OpenStudy (anonymous):

now you have \[x-4=3x\] and this should be more or less easy to solve

OpenStudy (anonymous):

idk how to solve that :/

OpenStudy (anonymous):

\[\frac{ 2 }{ 3\left( x-4 \right)}=2x\] 2x*3(x-4)=2 \[6x ^{2}-24x=2\] \[6x ^{2}-24x-2=0,3x ^{2}-12x-1=0\] \[x=\frac{ 12\pm \sqrt{\left(- 12 \right)^{2}-4*3*-1} }{2*3 }\] \[x=\frac{ 12\pm \sqrt{144+12} }{6 }=\frac{ 12\pm \sqrt{156} }{ 6 }\] now you can solve